Consider a vector of magnitude 3 units and in the direction of +X axis.This is followed by another vector 4 units parallel to the +y axis. What is the resultant of the two vectors


1
Expert's answer
2021-01-10T18:09:43-0500

"\sqrt{3^2+4^2}=5(units)"


"\tan\alpha=4/3\to \alpha=53.1°" from the positive direction of the x-axis